Jim cashed a paycheck and gave his son $10. He then spent $30 on fuel for his car and half of the remaining money on a new tent. He bought a pair of running shoes for $29.45 and had $17.75 left. How much did he receive when he cashed his paycheck?
Answer by richwmiller(17219)

(p-10-30) after son and gas
(p-10-30)/2 for tent
shoes 29.45
(p-40)-(p-40)/2-29.45=17.75
p-40-29.45-(p-40)/2=17.75
p-(p-40)/2=17.75+29.45
p-(p-40)/2=47.20
2p-p-40=94.40
p-40=94.40
p=134.40
check
134.40-10-30=94.40
94.40/2=47.20
47.20-29.45=17.75
ok